# PSA-MIL: A Probabilistic Spatial Attention-Based Multiple Instance Learning for Whole Slide Image Classification

### v1

@@ Abstract
Whole Slide Images (WSIs) are high-resolution digital scans widely used in medical diagnostics. WSI classification is typically approached using Multiple Instance Learning (MIL), where the slide is partitioned into tiles treated as interconnected instances. While attention-based MIL methods aim to identify the most informative tiles, they often fail to fully exploit the spatial relationships among them, potentially overlooking intricate tissue structures crucial for accurate diagnosis. To address this limitation, we propose Probabilistic Spatial Attention MIL (PSA-MIL), a novel attention-based MIL framework that integrates spatial context into the attention mechanism through learnable distance-decayed priors, formulated within a probabilistic interpretation of self-attention as a posterior distribution. This formulation enables a dynamic inference of spatial relationships during training, eliminating the need for predefined assumptions often imposed by previous approaches. Additionally, we suggest a spatial pruning strategy for the posterior, effectively reducing self-attention’s quadratic complexity. To further enhance spatial modeling, we introduce a diversity loss that encourages variation among attention heads, ensuring each captures distinct spatial representations. Together, PSA-MIL enables a more data-driven and adaptive integration of spatial context, moving beyond predefined constraints. We achieve state-of-the-art performance across both contextual and non-contextual baselines, while significantly reducing computational costs. Our code is available at https://github.com/SharonPeled/PSA-MIL.
Figure 1: Conceptual Illustration of Implied Spatial Relationships from an Anchored Patch in Different MIL Paradigms.
(a) Permutation-invariant MIL treats all patches equally, disregarding spatial relationships. (b) Contextual MIL restricts interactions to a predefined neighborhood around the anchor patch, enforcing a fixed spatial structure. (c) Our proposed PSA-MIL dynamically learns spatial relationships, starting with an initial distribution over neighboring patches and evolving throughout training.
@@ 1 Introduction
Histopathological image analysis plays a crucial role in modern medicine, often regarded as ground truth in medical practice [22, 15, 22, 28]. The digitization of pathological slides into Whole Slide Images (WSIs) has revolutionized the field of pathology, enabling large-scale computational analysis. Due to the immense size of a WSI, which can exceed a billion pixels, a widely adopted approach for computational analysis is to decompose these large images into smaller, processable patches. Exhibiting only slide-level annotations, histopathological image analysis is commonly formulated as a Multiple Instance Learning (MIL) task [42, 18], where each slide is modeled as a permutation-invariant bag of instances, overlooking spatial dependencies among instances.
@@
The incorporation of spatial context into MIL has gained popularity in recent WSI studies [13, 25, 9, 45]. While contextual models have shown strong performance, they often struggle to fully capture spatial dependencies among instances due to fixed spatial assumptions. For example, CAMIL [13] imposes predefined spatial constraints via a neighbor-constrained attention module, whereas Bayes-MIL [9] enforces spatial smoothness using a CRF with a fixed kernel size, both limiting flexibility in capturing diverse spatial patterns. Fig. 1 illustrates how different MIL paradigms model spatial relationships. Another popular approach for incorporating spatial context is positional encoding and its many variants [45, 25, 6, 36]. For example, SETMIL [45] employs Relative Positional Encoding (RPE) to model relative positions, HIPT [6] uses 2D absolute positional embeddings, and TransMIL [36] leverages PPEG, a convolution-based implicit positional encoding. While these methods have shown some performance gains, our experiments indicate that these improvements are often marginal. Moreover, their integration of spatial context is primarily empirically justified rather than theoretically grounded, resulting in a lack of a formal interpretative framework.
@@
To this end, we propose Probabilistic Spatial Attention MIL (PSA-MIL), a novel attention-based MIL approach that seamlessly integrates spatial context into the attention mechanism. Our method builds upon a probabilistic interpretation of self-attention, incorporating spatial relationships as learnable distance-decayed priors within a posterior distribution. This incorporation enables the model to capture how instances attend to each other based on their spatial proximity, balancing adaptive priors with observed likelihood. The main contributions of this paper can be summarized as follows:
@@
  • A theoretically grounded formulation for spatial modeling: We introduce a probabilistic self-attention framework that dynamically infers spatial dependencies during training rather than relying on predefined assumptions. This provides a principled and data-driven solution for leveraging spatial information in WSI analysis.
  • A spatial pruning strategy for computational efficiency: We propose a spatial pruning strategy for the posterior distribution, significantly reducing the quadratic complexity of self-attention while preserving critical spatial dependencies. This results in substantial computational savings without compromising performance.
@@
  • An entropy-based diversity loss for multi-head attention: We introduce a loss term that explicitly encourages spatial variation among attention heads, ensuring that each head captures distinct spatial representations. This mitigates redundant representations often observed in multi-head architectures [29, 43, 2].
Together, these contributions make PSA-MIL a more data-oriented and adaptive framework for integrating spatial context, moving beyond predefined constraints and enabling a principled approach to WSI analysis.
@@ 2.1 Multiple Instance Learning in WSI Analysis
Multiple Instance Learning (MIL) has become a widely adopted approach for WSI classification, allowing models to make slide-level predictions using only weakly supervised, slide-level annotations [4, 9, 42, 3]. Given the high resolution and large scale of WSIs, MIL frameworks process slides as bags of image tiles (instances), enabling models to learn relevant diagnostic patterns without requiring exhaustive manual annotations. Attention-based Multi-Instance Learning (ABMIL) was introduced, allowing models to learn instance-level importance weights [18]. Several works have since refined feature aggregation. For example, CLAM [27] incorporated a prototype-based clustering mechanism, and DTFD-MIL [42] employed a multi-tier fusion strategy. More recently, Transformer-based MIL architectures have emerged, leveraging self-attention to capture long-range dependencies between tiles [36, 13, 4, 45]. For example, TransMIL [36] utilized pairwise token interactions, enabling the model to encode global and local features. GTP [46] combined Graph Neural Networks (GNNs) with Transformers, enabling topology-aware modeling of spatial relationships within WSIs.
@@ 2.2 Spatial Context in Multiple Instance Learning
Spatial understandings of tissue structures are crucial during histopathological examination [35, 23, 17]. Reflecting this clinical importance, several approaches have been proposed for modeling spatial context in MIL. Multi-scale approaches, including HIPT [6], DS-MIL [24], and SETMIL [45], learn hierarchical representations spanning different magnifications. However, their hierarchical nature increases complexity, and adjacent tiles may be assigned to different regions, disrupting spatial continuity. Graph-based approaches explicitly model spatial relationships by treating tiles as graph nodes with structured dependencies [4, 46, 3]. For example, SM-MIL [4] applies a smoothness operator, enforcing gradual transitions in attention scores to better capture local dependencies. Although effective, these methods prune to over-smoothing, limiting the model’s ability to capture fine-grained spatial distinctions [33, 5]. Positional encoding methods introduce spatial bias into tile representations to enhance spatial awareness [25, 45, 6, 36]. For example, TransMIL [36] employs PPEG, a convolution-based positional encoding, while SETMIL [45] utilizes Relative Positional Encoding (RPE) to model spatial relationships. While proven effective, our experiments show that their performance gains are often marginal. Bayes-MIL [9] takes a different approach by incorporating spatial context through a Convolutional CRF, however, it still relies on a fixed kernel size, which restricts its ability to adapt. More recent approaches utilize spatially constrained Transformer-based MIL architectures, where interactions are restricted to fixed neighborhoods. For example, CAMIL [13] enforces predefined spatial constraints through a neighbor-constrained attention mechanism, while Li et al. [25] applies a fixed attention mask to model local interactions, both limiting flexibility in capturing diverse spatial patterns. Moreover, due to the quadratic complexity of self-attention, Transformer-based MIL architectures [36, 13] often rely on self-attention approximations, such as Nyströmformer [41], which can lead to suboptimal performance [10, 14]. In contrast, our PSA-MIL approach dynamically infers spatial relationships during training, adapting to the underlying tissue structure without rigid constraints. This paradigm shift is illustrated in Fig. 1. Additionally, our spatial pruning strategy eliminates the need for self-attention approximations, ensuring efficiency without compromising performance.
@@
Figure 2: PSA-MIL Overview: 1-2. Tissue regions are cropped from the WSI, and the resulting tiles are encoded using a pretrained feature extractor, producing tile representations. 3-3.1. A Multi-Head Spatial Self-Attention mechanism is employed to generate informative, spatially correlated tile representations. Within this block, the pairwise distance matrix (bottom left) undergoes dynamic spatial pruning via the learned function $f^{-1}(\tau|\theta)$, producing a spatially pruned distance matrix. The distance-decayed matrix is then computed using $f(d|\theta)$. Both $f^{-1}(\tau|\theta)$ and $f(d|\theta)$ are learnable components parameterized by $\theta$, enabling adaptive pruning and decay. Simultaneously, query, key, and value matrices are generated. The posterior attention distribution is then computed using Eq. 6 and applied to the value matrix, producing refined tile embeddings that encode spatial interactions. 4. These embeddings are subsequently pooled via an attention-based aggregation mechanism and passed through a classification head to generate slide-level predictions. The final loss function consists of both the classification loss and a diversity loss applied to the multi-head attention mechanism to encourage the capture of diverse spatial patterns.
@@ 3.1.1 Self-Attention for Multiple Instance Learning
Under the MIL framework, self-attention is utilized to model interactions among instances within a bag, enabling each instance (or tile) to attend to all other instances. Given a bag of instance embeddings $X\in\mathbb{R}^{n\times d}$, the self-attention operation computes queries $Q$, keys $K$, and values $V$ matrices through linear projections of the embeddings. The attention is defined as:
@@
where $H\in\mathbb{R}^{n\times d_{v}}$ represents the attention-weighted interactions among instances. Here, $d_{v}$ and $d_{k}$ denote the dimensionalities of the values and keys matrices, respectively.
Multi-head Attention Each output sequence $H$ forms an attention head. In multi-head attention, multiple heads are concatenated and then linearly projected to compute the final output:
[…]
where $W\in\mathbb{R}^{hd_{v}\times hd_{v}}$ is the projection matrix, and $h$ denotes the number of heads.
@@
While this formulation effectively captures intra-bag dependencies, it inherently treats the instances as unordered elements without considering their spatial relationships. This lack of spatial context may hinder the model’s ability to recognize more intricate structures.
@@ 3.1.2 Self-Attention for Multiple Instance Learning as Posterior Distribution
3.2 PSA-MIL for WSI Classification
In the context of WSIs, instances represent contiguous regions of tissue, and their spatial arrangement provides valuable diagnostic cues [35, 23, 17]. However, the self-attention MIL, as formulated in Sec. 3.1.2, inherently disregards spatial relationships among instances, which may limit the model’s ability to generalize. To this end, we propose PSA-MIL, a novel extension of the self-attention MIL that seamlessly integrates spatial context into the attention mechanism. Unlike previous approaches, our formulation enables an adaptive and principled spatial modeling without imposing predefined spatial constraints [13, 25, 9]. Building upon the probabilistic interpretation of self-attention described in Sec. 3.1.2, we relax both Assumption 1 ($\ell_{2}$-normalized queries and keys) and Assumption 2 (uniform priors), allowing for a more refined posterior estimation.
Relaxing Assumption 1 $\ell_{2}$-normalization of queries and keys is often introduced to align attention with a probabilistic interpretation [31]. However, standard self-attention does not inherently enforce this assumption, leading to inconsistencies in the formulation. We relax this assumption altogether and instead compute the full $\ell_{2}$-norm $||q_{i}-k_{j}||^{2}$, allowing queries and keys to retain magnitude information, which can enhance expressiveness [31].
@@ 3.2 PSA-MIL for WSI Classification
Pooling Operator The [CLS] token is a common pooling strategy in transformers [12], effective for aggregating information across multiple self-attention layers. Our model employs a single-layer, multi-head self-attention architecture for an optimal balance between performance and efficiency. Consequently, we adopt attention-based pooling instead of the [CLS] token, leveraging its enhanced interpretability.
@@ 3.2.1 Refined Posterior After Assumption Relaxation
${\textstyle p(t_{j}\!=\!1\!\mid\!q_{i})\!=\!\frac{\exp\left(\dfrac{\|q_{i}-k_{j}\|^{2}}{\sqrt{d_{k}}}+\log(f(d_{ij}|\theta))\right)}{\sum_{j^{\prime}=1}^{N}\!\exp\left(\dfrac{\|q_{i}-k_{j^{\prime}}\|^{2}}{\sqrt{d_{k}}}+\log(f(d_{ij^{\prime}}|\theta))\right)}}$ (6)
@@ 3.2.2 Parametric Distance-Decay Functions for Spatial Modeling
These parametric decay functions provide a structured, learnable mechanism for adjusting spatial influence, enabling the model to adaptively capture varying spatial dependencies.
Figure 3: The Effect of Diversity Loss on Multi-Head Attention. (a-b). Locality evolution: Without diversity loss, all heads converge to similar locality values, whereas with diversity loss, the locality values diverge. (c-d). Inter-head similarity: The similarity metric measures average $\ell_{2}$ token similarity (smoothed) across heads. Without diversity loss, similarity remains relatively high, while with diversity loss, similarity decreases, promoting diverse attention patterns. See Sec. 4.5 for additional details.
@@
Table 1: Performance comparison of different MIL approaches on TCGA-CRC and TCGA-STAD datasets, categorized into Non-Contextual and Contextual methods.
3.3 Spatially Pruning the Posterior
The quadratic computational complexity of self-attention presents a major challenge for long sequences, such as in WSI analysis. Proposed solutions include low-rank and kernel-based approximations [40, 8, 20, 41], such as Nyströmformer [41] used in TransMIL [36] and in CAMIL [13]. Though, these approaches often yield sub-optimal performance [10, 14]. Other methods apply local attention, limiting focus to local context while stacking multiple layers to capture global interactions [25]. Although local self-attention reduces computation, it ultimately relies on fixed spatial assumptions, which limits its ability to fully model spatial dependencies.
To address this challenge, we propose a spatial pruning strategy for the posterior distribution, leveraging the inferred prior distribution during training to restrict attention computations based on learned spatial dependencies. Namely, we apply a filtering criterion $f(d|\theta)\geq\tau$, where $\tau$ is a threshold parameter, permitting only instances with decays above $\tau$ to contribute to the posterior computation. Such pruning techniques have been explored in mixture models, demonstrating the effectiveness of removing components that contribute negligibly to the posterior [37, 32]. Notably, by utilizing an invertible parametric function $f(d|\theta)$, we establish that $d\leq f^{-1}(\tau|\theta)$, effectively reducing our method to a form of $K$-local self-attention with $K=\lceil f^{-1}(\tau|\theta)\rceil$ and a computational complexity of $O(n\cdot K^{2})\ll O(n^{2})$. Because $f(d|\theta)$ is learned during training, the locality of the self-attention is also learned, resulting in a dynamic local attention. Furthermore, unlike standard local self-attention, the retained instances are weighted by their learned decay values, making the operation spatially aware. Notably, we do not fine-tune $\tau$; instead, we typically set it to a constant value (e.g., $1e{-3}$), which we find works well in practice.
Fig. 2 illustrates how $f^{-1}(\tau|\theta)$ is applied in the PSA-MIL pipeline. Moreover, Fig. 3 (a-b) shows the evolution of our dynamic self-attention during training.

### v2

@@ Abstract
Whole Slide Images (WSIs) are high-resolution digital scans widely used in medical diagnostics. Due to their immense size, WSI classification is typically approached using Multiple Instance Learning (MIL), where a slide is partitioned into individual tiles, disrupting its spatial structure. Recent MIL methods often incorporate spatial context through rigid spatial assumptions (e.g. fixed kernels), which limit their ability to capture the intricate tissue structures crucial for an accurate diagnosis. To address this limitation, we propose Probabilistic Spatial Attention MIL (PSA-MIL), a novel attention-based MIL framework that integrates spatial context into the attention mechanism through learnable distance-decayed priors, formulated within a probabilistic interpretation of self-attention as a posterior distribution. This formulation enables a dynamic inference of spatial relationships during training, eliminating the need for predefined assumptions often imposed by previous approaches. Furthermore, we introduce a diversity loss that promotes complementary spatial representations across attention heads and a spatial posterior-pruning strategy that reduces computational cost for long WSI sequences while preserving performance. Extensive experiments across multiple datasets and tasks show that PSA-MIL outperforms current baselines and achieves state-of-the-art results with substantially lower computational overhead. Our code is available at https://github.com/SharonPeled/PSA-MIL.
Figure 1: Comparison of spatial context modeling approaches in WSI analysis. (a) Permutation-invariant MIL treats all patches equally, disregarding spatial relationships. (b) Contextual MIL restricts interactions to a predefined neighborhood around the anchor patch, enforcing a fixed spatial structure. (c) Positional encoding incorporates spatial bias into tile representations, providing limited spatial awareness. Its unstructured nature lacks alignment with physical distances, which hurts interpretability. (d) Our proposed PSA-MIL dynamically learns spatial relationships, starting with an initial distribution over neighboring patches and evolving throughout training, in contrast to previous methods that maintain a fixed spatial structure.
@@ 1 Introduction
Histopathological image analysis plays a crucial role in modern medicine, often regarded as ground truth in medical practice [21, 16]. The digitization of pathological slides into Whole Slide Images (WSIs) has revolutionized the field of pathology, enabling large-scale computational analysis. Due to the immense size of a WSI, which can exceed a billion pixels, a widely adopted approach for computational analysis is to decompose these large images into smaller, processable patches. Exhibiting only slide-level annotations, histopathological image analysis is commonly formulated as a Multiple Instance Learning (MIL) task [14], where each slide is modeled as a permutation-invariant bag of instances, overlooking spatial context among instances.
@@
Recognizing the critical role of spatial understanding of tissue structures during histopathological examination [34, 22, 18], recent advancements have introduced contextual MIL approaches which incorporate spatial relationships into the learning process [13, 4, 17, 11]. While contextual models have shown strong performance, they often struggle to fully capture spatial dependencies due to fixed spatial assumptions. For example, CAMIL [13] imposes predefined spatial constraints via a neighbor-constrained attention module, whereas Bayes-MIL [11] enforces spatial smoothness using a CRF with a fixed kernel size, both limiting flexibility in capturing diverse spatial patterns. Others, such as DAS-MIL[4] and H²-MIL[17], focus on distilling multi-scale and hierarchical representations. Another popular approach for incorporating spatial context is positional encoding and its many variants [44, 23, 7, 35]. For example, SETMIL [44] employs Relative Positional Encoding (RPE) to model relative positions, and TransMIL [35] leverages PPEG, a convolution-based implicit positional encoding. However, our experiments indicate that their improvements are often marginal, and their integration of spatial context is primarily empirically justified rather than theoretically grounded, resulting in a lack of a formal interpretative framework.
Furthermore, long sequences, such as those in WSI analysis, pose a challenge for transformer-based architectures due to the quadratic complexity of self-attention. Solutions like low-rank and kernel-based approximations, including Nyströmformer [39] used in TransMIL [35] and in CAMIL [13], often deliver suboptimal performance [12, 15]. Other methods use local attention to reduce computations [23], but are limited to fixed spatial context.
@@
To this end, we propose Probabilistic Spatial Attention MIL (PSA-MIL), a novel attention-based MIL approach that seamlessly integrates spatial context into the attention mechanism. Our method builds upon a probabilistic interpretation of self-attention, incorporating spatial relationships as learnable distance-decayed priors within a posterior distribution. This incorporation enables the model to capture how instances attend to each other based on their spatial proximity, balancing adaptive priors with observed likelihood. Fig. 1 compares existing spatial modeling approaches with our proposed PSA-MIL. The main contributions of this paper can be summarized as follows:
@@
  • A Principled Self-Attention Formulation for Adaptive Contextual MIL: We introduce a probabilistic self-attention framework that learns spatial dependencies dynamically during training. This data-driven approach enhances spatial context modeling in WSI analysis, overcoming the limitations of prior methods that rely on fixed assumptions and are often subject to extensive tuning.
  • Dynamic Multi-Head Spatial Specialization: Our formulation enables each attention head to independently derive its own spatial interaction range — a property unique to our approach.
@@
  • An Entropy-based Diversity Loss for Multi-head Attention: We introduce a loss term that explicitly encourages spatial variations among attention heads, ensuring that each head captures distinct spatial representations. This mitigates redundant representations often observed in multi-head architectures [27, 43, 2].
  • A Spatial Pruning Strategy for Computational Efficiency: We propose a spatial pruning strategy for the posterior estimation that significantly reduces the quadratic complexity of self-attention while preserving critical spatial dependencies. This approach achieves substantial computational savings without compromising performance, in contrast to other low-cost self-attention approximations that may yield suboptimal solutions [23, 12, 15].
@@ 2.1 Multiple Instance Learning in WSI Analysis
Multiple Instance Learning (MIL) has become a widely adopted approach for WSI classification, allowing models to make slide-level predictions from weakly supervised annotations by treating whole slides as bags of image tiles. [5, 11, 42, 3]. Attention-based Multi-Instance Learning (ABMIL) allowed to learn instance-level importance weights [19]. Several works have since refined feature aggregation. For example, CLAM [26] incorporated a prototype-based clustering mechanism, and DTFD-MIL [42] employed a multi-tier fusion strategy. More recently, Transformer-based MIL architectures have emerged, leveraging self-attention to capture long-range dependencies between tiles [35, 13, 5, 44]. For example, TransMIL [35] modeled global and local features via token interactions, while GTP [45] combined GNNs and Transformers for topology-aware spatial modeling.
@@ 2.2 Spatial Context in Multiple Instance Learning
Reflecting the importance of spatial understanding in histopathology [34, 22, 18], several approaches have been proposed for modeling spatial context in MIL. Graph-based approaches model spatial relationships by treating tiles as graph nodes with structured dependencies [5, 45, 3]. For example, SM-MIL [5] applies a smoothness operator, enforcing gradual transitions in attention scores to better capture local dependencies. Although effective, these methods are prune to over-smoothing, limiting the model’s ability to capture fine-grained spatial distinctions [31, 6]. Additional analysis and comparisons of GNN-based spatial modeling are provided in the supplementary material. Bayes-MIL [11] takes a different approach by incorporating spatial context through a Convolutional CRF, however, it still relies on a fixed kernel size, which restricts its ability to adapt.
@@
Another direction for modeling spatial context involves positional encoding, which embeds spatial bias into tile representations to enhance spatial awareness [23, 44, 7, 35]. For example, TransMIL [35] employs PPEG, a convolution-based positional encoding, while SETMIL [44] utilizes Relative Positional Encoding (RPE) to model spatial relationships. While proven effective, our experiments show that their performance gains are often marginal. More recent approaches utilize spatially constrained Transformer-based MIL architectures, where interactions are restricted to fixed neighborhoods. For example, CAMIL [13] enforces predefined spatial constraints through a neighbor-constrained attention mechanism, while LongMIL [23] applies a fixed attention mask to model local interactions, both limiting flexibility in capturing diverse spatial patterns. Moreover, due to the quadratic complexity of self-attention, Transformer-based MIL architectures [35, 13] often rely on self-attention approximations, such as Nyströmformer [39], which can lead to suboptimal performance [12, 15]. In contrast, our PSA-MIL approach dynamically infers spatial relationships during training, adapting to the underlying tissue structure without rigid constraints. This paradigm shift is illustrated in Fig. 1. Additionally, our spatial pruning strategy eliminates the need for self-attention approximations, ensuring efficiency without compromising performance.
@@
Figure 2: PSA-MIL Overview: 1–2. Tissue regions are cropped and encoded using a pretrained feature extractor, producing tile representations. 3-3.1. A multi-head spatial self-attention mechanism is employed to generate informative, spatially correlated tile representations. Within this block, the pairwise distance matrix (bottom left) undergoes dynamic spatial pruning (Sec. 3.4) via the learned function $f^{-1}(\tau|\theta)$, producing a spatially pruned distance matrix. The distance-decayed matrix is then computed using $f(d|\theta)$. Both $f^{-1}(\tau|\theta)$ and $f(d|\theta)$ are learnable components parameterized by $\theta$, enabling adaptive pruning and decay. Simultaneously, query, key, and value matrices are generated. The posterior attention distribution is then computed using Eq. 6 and applied to the value matrix, producing refined tile embeddings that encode spatial interactions. 4. These embeddings are subsequently pooled via an attention-based aggregation mechanism and passed through a classification head to generate slide-level predictions. The final loss function consists of both the classification loss and a diversity loss (Sec. 3.3) applied to the multi-head attention mechanism to encourage the capture of diverse spatial patterns.
@@ 3.1.1 Self-Attention for Multiple Instance Learning
Under the MIL framework, self-attention enables each instance (or tile) to interact with all other instances in the bag. The attention is computed as:
@@
where $Q$, $K$, and $V$ are linear projections of the instance embeddings and $d_{k}$ denotes the dimension of Keys.
Multi-head Attention: Multiple attention heads are concatenated and projected to produce the final output:
$\text{Multi-Head}(\{Q,K,V\}_{i=1}^{h})=\text{Concat}(H_{1},\ldots,H_{h})W$
@@
While this approach captures intra-bag dependencies, it ignores spatial relationships among instances, which may limit the model’s ability to capture intricate structures.
@@ 3.1.2 Self-Attention for Multiple Instance Learning as Posterior Distribution
3.2 Learnable Probabilistic Spatial Attention for Adaptive Contextual MIL
The self-attention MIL, as formulated in Sec. 3.1.2, inherently disregards spatial relationships among instances, which is crucial for WSI analysis, where instances represent contiguous regions of tissue and their spatial arrangement provides valuable diagnostic cues [34, 22, 18]. To this end, we propose PSA-MIL, an innovative approach that seamlessly integrates spatial context into the attention mechanism as a learnable component, enabling the model to capture spatial dependencies in a flexible and data-driven manner. Building upon the probabilistic interpretation of self-attention described in Sec. 3.1.2, we relax both Assumption 1 ($\ell_{2}$-normalized queries and keys) and Assumption 2 (uniform priors), allowing for a more refined posterior estimation.
Relaxing Assumption 1 $\ell_{2}$-normalization of queries and keys is often introduced to align attention with a probabilistic interpretation [29]. However, standard self-attention does not inherently enforce this assumption, leading to inconsistencies in the formulation. We relax this assumption altogether and instead compute the full $\ell_{2}$-norm $q_{i}-k_{j}||^{2}$, allowing queries and keys to retain magnitude information, which can enhance expressiveness [29].
@@ 3.2 Learnable Probabilistic Spatial Attention for Adaptive Contextual MIL
Pooling Operator We adopt attention-based pooling due to its effectiveness in aggregating instance-level features and its interpretability. This is a common strategy in WSI analysis [13, 5]. Additional details regarding our architecture design are provided in the supplementary material.
@@ 3.2.1 Refined Posterior After Assumption Relaxation
$p(t_{j}\!=\!1\!\mid\!q_{i})\!=\!\frac{\exp\left(\dfrac{\|q_{i}-k_{j}\|^{2}}{\sqrt{d_{k}}}+\log(f(d_{ij}|\theta))\right)}{\sum_{j^{\prime}=1}^{N}\!\exp\left(\dfrac{\|q_{i}-k_{j^{\prime}}\|^{2}}{\sqrt{d_{k}}}+\log(f(d_{ij^{\prime}}|\theta))\right)}$} (6)
@@ 3.2.2 Parametric Distance-Decay Functions for Spatial Modeling
These parametric decay functions provide a structured, learnable mechanism for modeling spatial influence. This represents a paradigm shift, enabling principled spatial modeling without imposing predefined constraints, as seen in previous approaches [13, 11]. See Fig. 1 for illustration.
Figure 3: The Effect of Diversity Loss on Multi-Head Attention. (a-b). Locality evolution: In Sec. 3.4, we describe how our spatial attention can be interpreted as dynamic local attention. Without diversity loss, all heads converge to similar locality values, whereas with diversity loss, the locality values diverge. (c-d). Inter-head similarity: The similarity metric measures average $\ell_{2}$ token similarity (smoothed) across heads. Without diversity loss, similarity remains relatively high, while with diversity loss, similarity decreases. See Sec. 4.4.2 for additional details.
@@
Table 1: Subtyping performance of different MIL approaches on TCGA-CRC and TCGA-STAD datasets, categorized into Non-Contextual and Contextual methods. PSA-MIL variants achieve top performance across both tasks, demonstrating the benefits of our spatial modeling.